OpenStudy (anonymous):

Find the rectangular coordinates of the point with the given polar coordinates. (-4,-pi/3)

OpenStudy (rogue):

Polar coordinates:\[(r , \theta)\]Rectangular:\[(x, y)\]Polar to rectangular:\[(r, \theta) \rightarrow (r \cos \theta, r \sin \theta)\]

OpenStudy (rogue):

Can you do the rest?

OpenStudy (anonymous):

yeah i already got the answer. thank you though:)

OpenStudy (rogue):

Alright, no problem =)